OpenStudy (anonymous):

Find the distance between the lines with equations -9x-8y+4=0 and y=-9/8x+6

OpenStudy (anonymous):

well,how about finding a point of this equation? -9x - 8y + 4 = 0 --> 8y = -9x + 4 --> y = -9/8 x + 1/2

OpenStudy (anonymous):

can you find a point?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

can you show me the formula and how to do it please

OpenStudy (anonymous):

ok,i will show you,but 8 is not correct,i said find a point!and points are used to write like this: (x,y)

OpenStudy (anonymous):

oh okay. well im trying to not graph it

OpenStudy (anonymous):

(8,-8.5) is a point of this line...

OpenStudy (anonymous):

-8.5 =( -9/8 x 8) + 1/2

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

howd you get (8,-8.5) ?

OpenStudy (anonymous):

it's not difficult,you just need a table like this:

OpenStudy (anonymous):



OpenStudy (anonymous):

for example if x=1 then find y from this equation: y = (-9/8 x 1) + 1/2

OpenStudy (anonymous):

i almost have to go :( whats the answer you got?

OpenStudy (anonymous):

just keep typing me all the things you did to get the answer

OpenStudy (anonymous):

i didn't calculate,but i think it's much better that you know how to solve it,if you get answer,you can't solve another same questions,can you?

OpenStudy (anonymous):

we have a formula for these questions...but i wanted to tell you how to solve it,but if you want the formula,i will give it to you :)

OpenStudy (anonymous):

no i cant.. but i just wanted to see how you work it out and the answer so i could see exactly how you did it in order for me to try another problem like that

OpenStudy (anonymous):

ii gotta go :(

OpenStudy (anonymous):

ok when we have these two equations: y = mx + b1 y = mx + b2 use this formula: \[d = \frac{ |b2 - b1| }{ \sqrt{m ^{2}+1} }\]

OpenStudy (anonymous):

but for solving it,we can use this formula or solve it by the laws of mathematics...

OpenStudy (anonymous):

i can tell you how to solve it by the laws of mathematics :)